# Observation of Power Superbroadening of Spectral Line Profiles on IBM Quantum

### v2

@@ III The quadratic and the even exponent pulse shapes
Based on the sought properties in the superadiabatic basis, two families of temporal pulse shapes are introduced—the quadratic and the even exponent pulse shapes. The quadratic pulse shape resembles a rectangular pulse with an added parabolic term, as illustrated in Fig. 2(a). Despite the simplicity of its crude form – $\Omega(t)\propto 1+\beta t^{2}$, where $\beta$ is a parameter – the normalized expression for this pulse shape is
@@
Similarly designed to break the superadiabatic inequality, the even-exponent pulse shape incorporates nonadiabatic features, namely discontinuous jumps at the edges and an inherent zero value at the midpoint of the pulse. Its crude form is $\Omega(t)\propto t^{2P}$, but adjusted for unity it becomes
@@
where $P\in\mathbb{N}$ is a parameter, half the exponent of time $t$, and controls the width of the pit. With $P=0$, the even exponent (EE) pulse shape becomes the standard rectangular pulse shape, or put alternatively – the pit has zero width. For $P\gg 1$ the shape approximates two Dirac delta functions time $T$ apart. For a high exponent $P$ and the right pulse area, where each of the two horns has area $\mathcal{A}=\pi/2$, combining for a total area of $\pi$, the pulse can be used to produce Ramsey fringes in the frequency spectrum. Similar to the $\beta=1$ case of the quadratic pulse shape, the even-exponent pulse also inherently reaches zero in the midpoint for all values of its parameter $P$.
Interestingly, the $\beta=-1$ and $\beta=+1$ quadratic pulse shapes have the same derivatives in magnitude, only the signs are inverted. Intuitively, this could incline us to believe that their representations in the adiabatic basis would not differ and that the adiabatic inequality would lead to similar conclusions for both of them. However, one can see in Fig. 1 that the superadiabatic inequality is violated by the $\beta=1$ quadratic pulse, and by the $N=3$ even-power pulse shape but not by the $\beta=-1$ quadratic pulse shape. This finally allows us to see why the $\beta=-1$ and $\beta=+1$ shapes behave differently at $\Delta\neq 0$.
@@ III.1 Orientation-changing quadratic
The orientation-changing quadratic pulse shapes with $\beta=-1,0,0.25,0.5,0.75\text{ and }1$ were tested on the 127-qubit IBM quantum computer – ibm_sherbrooke. The detailed list of parameters recorded during the experiment is shown in Sec. A. The duration of all orientation-changing quadratic pulses was 800 dt units—the system time units used in IBM’s Qiskit framework—which corresponds to 177.8 ns. As stated previously, the $\beta=0$ case of the face-changing quadratic is in fact the constant pulse. It arises from simply turning the pulse generator on and off after time $T$ and is known for reproducing the common linear power broadening. In line with this hypothesis, we indeed see the common broadening pattern in the top right excitation map of Fig. 2(b), corresponding to the $\beta=0$ case. It is also evident that as $\beta$ changes from -1 to 0, the power broadening pattern grows stronger. In the $\beta=-1$ plot the second maximum is about $20$ MHz wide, whereas in the $\beta=0$ case, a fringe appears at around $\Delta=\pm 15$ MHz, resulting in a linewidth of 30 MHz. The broadening pattern is much stronger in the linewidth of the $\beta=1$ pulse, spanning a frequency interval of about 100 MHz, or an increase by a factor of about 3.3 compared to the rectangular pulse ($\beta=0$). It is not intuitive that the $\beta=-1$ case, albeit having the same derivative in magnitude, shows strikingly weaker power broadening than the $\beta=1$ case. These two examples form the limiting cases in the experiment, with the $\beta=1$ plot having at least 5 visible fringes around the second maximum. Both pulse shapes have exactly the same derivatives in magnitude, but their signs are different. In the $\beta=-1$ case, the derivative is first positive, then negative for the second half of the pulse, while in the $\beta=1$ case, they are in the opposite order. This difference may seem insignificant at first, but completely changes the pulse’s appearance in the adiabatic and superadiabatic pictures.
@@
III.2 Even-exponent pulse
The excitation landscapes in Fig. 3(b) demonstrate progressively stronger power broadening, accompanied by increasingly pronounced fringe patterns. Aiming at higher non-adiabaticity, we evaluated even-exponent pulse shapes with $P\in\{0,1,2,3\}$ on the 127-qubit IBM quantum computer – ibm_sherbrooke. The values of key parameters recorded throughout the demonstration are listed in Sec. A. The duration of all even-exponent pulses was 1600 dt units or 355.6 ns. Notably, for all $P>0$, the Rabi frequency at the pulse midpoint is zero, a distinctive feature of the even-exponent pulse shape. This characteristic enhances the pulse’s nonadiabatic behavior, effectively mimicking a configuration resembling two well-separated pulses, each with an area of $\pi/2$, similar to those employed in Ramsey interferometry to generate Ramsey fringes. Taking a look at the width of the spectral lines for a $\pi$-area pulse for $P=0$ and $P=3$, we see a huge broadening in the $P=3$ case. A glance at the first maxima standing at $\mathcal{A}=\pi$ signals that it broadens from roughly $8$ MHz wide in the $P=0$ case to about $30$ MHz in the $P=3$ case. While the linewidth of the $3\pi$ pulse is about $20$ MHz in the $P=0$ case, it is wider than the $70$ MHz scope of the plot depicting the $P=3$ case, obtained by taking all visible fringes into account.
Examining the two plots on the bottom row — corresponding to $P=2$ (left) and $3$ (right) — we find that dark nearly vertical lines have started to appear $\approx 3$ MHz apart due to interference between the left and right horns of the even-exponent pulse. These lines are to increase both in number and contrast with higher numbers of $P$.
@@ V Acknowledgments
Appendix A Quantum processor specs
@@
Due to the limited free slots on the hardware, the measurements were taken on 5 different days. The $P=0$ even-exponent pulse was taken on 27 Nov 2024, while the $P=1,2,3$ pulses were recorded on 29 Nov 2024. The $\beta=-1$ and 00 quadratic pulses were demonstrated on 1 Dec 2024, the $\beta=1$ pulse test was performed on 3 Dec 2024, and the $\beta=0.75,0.5$ and $0.25$ quadratic pulses — on 5 Dec 2024.
Qubit 46 of the ibm_sherbrooke system had a excitation frequency of 4.6741 GHz and an anharmonicity of -0.3134 GHz, both of which were constant over the days of the demonstration. The $T_{1}$ decoherence times were between 250 and 400 µs, while the $T_{2}$ were even better — between 280 and 700 µs. A more detailed breakdown of the parameters on each day of the demonstration can be found in Table 1. The recorded values imply a $T_{2}$-limited linewidth of several kHz, which is tiny compared to a linewidth between 5 and 150 MHz. Thus, the dephasing effects in this demonstration are insignificant.

### v3

@@ III The quadratic and the power-law pulse shapes
The quadratic pulse shape can be regarded as a rectangular pulse with an added parabolic term, as illustrated in Fig. 2(a). Despite the simplicity of its crude form – $\Omega(t)\propto 1+\beta t^{2}$, where $\beta$ is a parameter – the normalized expression for this pulse shape is
@@
Similarly designed to break the superadiabatic inequality, the power-law pulse shape incorporates nonadiabatic features, namely discontinuous jumps at the edges and an inherent zero value at the midpoint of the pulse. Its crude form is $\Omega(t)\propto t^{2P}$, but adjusted for unity it becomes
@@
where $P\in\mathbb{N}$ is a parameter, half the exponent of time $t$, and controls the width of the pit. With $P=0$, the power-law pulse shape becomes the standard rectangular pulse shape, or put alternatively – the pit has zero width. For $P\gg 1$ the shape approximates two Dirac delta functions time $T$ apart. For a high power $P$ and the right pulse area, where each of the two horns has an area $\mathcal{A}=\pi/2$, combining for a total area of $\pi$, the pulse can be used to produce Ramsey fringes in the frequency spectrum. Similar to the $\beta=1$ case of the quadratic pulse shape, the power-law pulse also inherently reaches zero in the midpoint for all values of its parameter $P$.
Interestingly, the $\beta=-1$ and $\beta=+1$ orientation-changing quadratic pulse shapes have the same derivatives in magnitude, only the signs are inverted. Intuitively, this could incline us to believe that their representations in the adiabatic basis would not differ and that the adiabatic inequality would lead to similar conclusions for both of them. However, one can see in Fig. 1 that the superadiabatic inequality is violated by the $\beta=1$ quadratic pulse, and by the $P=3$ power-law pulse shape but not by the $\beta=-1$ quadratic pulse shape. This finally allows us to see why the $\beta=-1$ and $\beta=+1$ shapes behave differently at $\Delta\neq 0$.
@@ III.1 Orientation-changing quadratic
The orientation-changing quadratic pulse shapes with $\beta=-1,0,0.25,0.5,0.75\text{, and }1$ were tested on the 127-qubit IBM quantum computer — ibm_sherbrooke. The detailed list of parameters recorded during the experiment is shown in Appendix A. The duration of all quadratic pulses was 800 dt units — the system time units ($dt=\frac{2}{9}$ ns) used in IBM’s Qiskit framework — which corresponds to 177.8 ns. As stated previously, the $\beta=0$ case of the orientation-changing quadratic is in fact the constant pulse. It arises from simply turning the pulse generator on and off after time $T$ and is known for reproducing the common linear power broadening. In line with this hypothesis, we indeed see the common broadening pattern in the top right excitation map of Fig. 2(b), corresponding to the $\beta=0$ case. It is also evident that as $\beta$ changes from -1 to 0, the power broadening pattern grows stronger. In the $\beta=-1$ plot the second maximum is about $20$ MHz wide, whereas in the $\beta=0$ case, a fringe appears at around $\Delta=\pm 15$ MHz, resulting in a linewidth of 30 MHz. The broadening pattern is much stronger in the linewidth of the $\beta=1$ pulse, spanning a frequency interval of about 100 MHz, or an increase by a factor of about 3.3 compared to the rectangular pulse ($\beta=0$). It is not intuitive that the $\beta=-1$ case, albeit having the same derivative in magnitude, shows strikingly weaker power broadening than the $\beta=1$ case. These two examples form the limiting cases in the experiment, with the $\beta=1$ plot having at least 5 visible fringes around the second maximum. Both pulse shapes have exactly the same derivatives in magnitude, but their signs are different. In the $\beta=-1$ case, the derivative is first positive, then negative for the second half of the pulse, while in the $\beta=1$ case, they are in the opposite order. This difference may seem insignificant at first, but completely changes the pulse’s appearance in the adiabatic and superadiabatic pictures.
@@
III.2 Power-law pulse
The excitation landscapes in Fig. 3(b) demonstrate progressively stronger power broadening, accompanied by increasingly pronounced fringe patterns. Aiming at higher non-adiabaticity, we evaluated power-law pulse shapes with $P\in\{0,1,2,3\}$ on the 127-qubit IBM quantum computer – ibm_sherbrooke. The values of key parameters recorded throughout the demonstration are listed in Appendix A. The duration of all power-law pulses was 1600 dt units, or 355.6 ns. Notably, for all $P>0$, the Rabi frequency at the pulse midpoint is zero, a distinctive feature of the power-law pulse shape. This characteristic enhances the pulse’s nonadiabatic behavior, effectively mimicking a configuration resembling two well-separated pulses, each with an area of $\pi/2$, similar to those employed in Ramsey interferometry to generate Ramsey fringes. Taking a look at the width of the spectral lines for a $\pi$-area pulse for $P=0$ and $P=3$, we see a huge broadening in the $P=3$ case. A glance at the first maxima standing at $\mathcal{A}=\pi$ signals that it broadens from roughly $8$ MHz wide in the $P=0$ case to about $30$ MHz in the $P=3$ case. While the linewidth of the $3\pi$ pulse is about $20$ MHz in the $P=0$ case, it is wider than the $70$ MHz scope of the plot depicting the $P=3$ case, obtained by taking all visible fringes into account.
Examining the two plots on the bottom row — corresponding to $P=2$ (left) and $3$ (right) — we find that dark nearly vertical lines have started to appear $\approx 3$ MHz apart due to interference between the left and right horns of the power-law pulse. These lines are to increase both in number and contrast with higher numbers of $P$.
@@ V Acknowledgments
Appendix A Quantum hardware demonstration specifics
@@
Due to the limited free slots on the hardware, the measurements were taken on 5 different days. The $P=0$ power-law pulse was taken on 27 Nov 2024, while the $P=1,2,3$ pulses were recorded on 29 Nov 2024. The $\beta=-1$ and 0 quadratic pulses were demonstrated on 1 Dec 2024, the $\beta=1$ pulse test was performed on 3 Dec 2024, and the $\beta=0.75,0.5$ and $0.25$ quadratic pulses — on 5 Dec 2024.
Qubit 46 of the ibm_sherbrooke system had a excitation frequency of 4.6741 GHz and an anharmonicity of -0.3134 GHz, both of which were constant over the days of the demonstration. The $T_{1}$ decoherence times were between 250 and 400 µs, while the $T_{2}$ were even better — between 280 and 700 µs. A more detailed breakdown of the parameters on each day of the demonstration can be found in Table 1. The recorded values imply a $T_{2}$-limited linewidth of several kHz, which is tiny compared to a linewidth between 5 and 150 MHz. Thus, the dephasing effects in this demonstration are insignificant. Each datapoint in the excitation landscapes of Figs. 2(b) and 3(b) is an average of 200 shots.